# Taming Self-Training for Open-Vocabulary Object Detection

### v2

@@ 5 Conclusion
In this paper, we highlight two challenges associated with applying self-training to OVD: 1) noisy PLs from pretrained VLMs and 2) frequent changes of PLs’ distributions. To address the challenges, we introduce a split-and-fusion (SAF) head and implement periodic updates. The SAF head splits a standard detection head into an open-branch and a closed-branch. The open-branch is trained with the class labels of ground truth and PLs. The closed-branch is trained with both class labels and boxes of ground truth. This approach effectively eliminates the location noise of PLs during training. Furthermore, the SAF head acquires complementary knowledge from different sets of training data, enabling fusion to enhance performance. The periodic updates reduce the number of updates to the teacher models, thereby decreasing the frequency of changes in PLs’ distributions. We demonstrate that the proposed method SAS-Det is both efficient and effective. Our pseudo labeling is much faster than prior PL-based methods [12, 57]. SAS-Det also outperforms recent models of the same scale on both COCO and LVIS for OVD.
@@ A.1 Our detector for OVD without an external RPN
Method | AP (COCO) | AP (LVIS) |
@@ A.2 Improving initial pseudo labels with RPN scores
Method | Training Data | Backbone | Detector | AP | AP | AP | AP |

### v3

@@ 5 Conclusion
In this paper, we highlight two challenges associated with applying self-training to OVD: 1) noisy PLs from pretrained VLMs and 2) frequent changes of PLs’ distributions. To address the challenges, we introduce a split-and-fusion (SAF) head and implement periodic updates. The SAF head splits a standard detection head into an open-branch and a closed-branch. The open-branch is trained with the class labels of ground truth and PLs. The closed-branch is trained with both class labels and boxes of ground truth. This approach effectively eliminates the location noise of PLs during training. Furthermore, the SAF head acquires complementary knowledge from different sets of training data, enabling fusion to enhance performance. The periodic updates reduce the number of updates to the teacher models, thereby decreasing the frequency of changes in PLs’ distributions. We demonstrate that the proposed method SAS-Det is both efficient and effective. Our pseudo labeling is much faster than prior PL-based methods [12, 59]. SAS-Det also outperforms recent models on both COCO and LVIS for OVD.
Acknowledgments: This research project has been partially funded by research grants to Dimitris N. Metaxas through NSF: 2310966, 2235405, 2212301, 2003874, and FA9550-23-1-0417.
@@ A.1 Our detector for OVD without an external RPN
Method | AP (COCO) | APr (LVIS) |
@@ A.2 Improving initial pseudo labels with RPN scores
Method | Training Data | Backbone | Detector | APr | APc | APf | AP |